# Evaluating and improving wave and non-wave stress parametrisations for oceanic flows

### v1

@@ 5.2 Suggested steady flow parametrisations
For a purely steady flow, we found that the parametrisations due to Bell (1975) and Klymak et al. (2010) had mostly accurate scalings, but required additional parameters to be accurate across a diverse set of regimes.
@@
where $A_{1}$, $B_{1}$ and $C_{1}$ are tunable constants, $Fr=Nh_{0}/U$, and $F_{\text{Bell2d}}$ is as defined in (2.2.1). Our data gave the fitted values $A_{1}=0.71$, $B_{1}=0.46$ and $C_{1}=1.4$. Note that one can use the Taylor expansion of (39) for small Froude numbers ($Fr\ll 1$) to obtain an expression with polynomial dependence on $Fr$.
@@
$F_{\text{3d}}$ $=\frac{1}{A_{2}-B_{2}|fW/U|}F_{\text{Bell3d}},\quad\text{if}\ Fr\lesssim 1,\ \text{and}\ |fW/U|\lesssim 1$ (41)
@@ 5.3 Suggested mixed flow parametrisations
$U(t)=U_{\text{mean}}+U_{\text{tidal}}\cos(\omega t)$ (43)
@@
[…]
@@
proved to be accurate for a wide range of parameters. Here, as in Sections 34.5 and 34.6, $\widetilde{F}_{m}$ and $\widetilde{F}_{p}$ are the values of $F_{m}$ and $F_{p}$ if $U_{\text{tidal}}$ or $U_{\text{mean}}$ are set to be zero respectively. In two dimensions, the main regimes where (44) disagreed with our simulations were those near the critical latitude $|f|=\omega$, those with small values of $U_{\text{mean}}$ (for $F_{m}$), and those with small values for $W$ (for $F_{p}$). Then, in three dimensions, the naive parametrisation was again less accurate near the critical latitude $|f|=\omega$, and also for large hill heights $h_{0}$ (for $F_{m}$). Further work is required to accurately parametrise the stress in these specific cases.
@@ 6 Conclusion
We have evaluated existing stress parametrisations for tidal, steady and mixed flows over rough topography. In the case where the topography is an isolated Gaussian hill in two or three dimensions, we have directly compared these parametrisations with idealised numerical simulations. In some regimes, such as those with large scale topography, we find that existing parametrisations fall short. As a result, we have suggested new or improved parametrisations that more accurately match the simulation data. It is hoped that these new parametrisations may be used to more accurately model stress for mixed flows or other cases for which there is currently limited theory available. In future work, it would be insightful to perform a similar analysis in less-idealised settings with more complicated topography.

### v3

@@ 5.2 Suggested steady flow parametrisations
For a purely steady flow, we found that the parametrisations due to Bell (1975) and Klymak et al. (2010) had mostly accurate scalings. However, to better match the simulation data we suggest introducing additional parameters that account for non-linear effects and need to be suitably chosen.
@@
where $A_{1}$, $B_{1}$ and $C_{1}$ are tunable constants, $Fr=Nh_{0}/U$, and $F_{\text{Bell2d}}$ is as defined in (2.2.1). Our data gave the fitted values $A_{1}=0.71$, $B_{1}=0.46$ and $C_{1}=1.4$. Note that one can use the Taylor expansion of (44) for small Froude numbers ($Fr\ll 1$) to obtain an expression with polynomial dependence on $Fr$. The extreme cases of low and high Froude numbers covers many possible regimes in the real world, for example abyssal hills for low $Fr$ and seamounts for large $Fr$. However, we did not analyse the transitional case of medium Froude numbers ($Fr\approx 1-10$) in detail. For this, we refer the reader to the recent work of Klymak et al. (2021) where a hybrid parametrisation is tested to account for this additional regime.
@@
$F_{\text{3d}}$ $=\frac{1}{A_{2}-B_{2}|fW/U|}F_{\text{Bell3d}},\ \text{if}\ Fr\lesssim 1,\ \text{and}\ |fW/U|\lesssim 1,$ (46)
@@ 5.3 Suggested mixed flow parametrisations
$U(t)=U_{\text{mean}}+U_{\text{tidal}}\cos(\omega t),$ (48)
@@
$F_{2d}$ $=F_{m}+F_{p}\cos(\omega t+\phi),\qquad\text{(two dimensions)}$
$F_{3d}$ $=F_{m}+F_{p}\cos(\omega t+\phi),\qquad\text{(three dimensions)}$
@@
proved to be accurate for a wide range of parameters. Here, as in Sections 44.5 and 44.6, $\widetilde{F}_{m}$ and $\widetilde{F}_{p}$ are the values of $F_{m}$ and $F_{p}$ if $U_{\text{tidal}}$ or $U_{\text{mean}}$ are set to be zero respectively. In two dimensions, the main regimes where (49) disagreed with, and notably underestimated, our simulation output were those near the critical latitude $|f|=\omega$, those with small values of $U_{\text{mean}}$ (for $F_{m}$), and those with small values for $W$ (for $F_{p}$). Then, in three dimensions, the naive parametrisation again underestimated the simulation data near the critical latitude $|f|=\omega$ (for $F_{p}$), and also for large hill heights $h_{0}$ (for $F_{m}$). For this last point, we suggest setting
$F_{m}=\widetilde{F}_{m}\left(1+\frac{U_{\text{tidal}}^{2}}{2U_{\text{mean}}^{2}}\right),$ (50)
in three dimensions whenever the Froude number $Fr=Nh_{0}/U_{\text{mean}}$} is large. The formula (50) is derived from the quadratic drag law (47) we observed in the steady flow case for $Fr\gg 1$. However, as discussed in Section 4f, it would be preferable to do more tests and analysis to evaluate the accuracy of (50), and to better understand the transition between the low and high Froude number regimes.
@@ 6 Conclusion
We have evaluated common stress parametrisations for tidal, steady and mixed flows over rough topography. In the case where the topography is an isolated Gaussian hill in two or three dimensions, we have directly compared these parametrisations with idealised numerical simulations. This helped clarify the cases for which the parametrisations and simulations agreed and disagreed. We found that the most notable regimes where the parametrisations and simulations disagreed were (i) when the topographic height $h_{0}$ was large, and (ii), when the magnitude of the Coriolis parameter $|f|$ was close to the tidal frequency $\omega$ (for tidal and mixed flows). This was not entirely unexpected, as the existing parametrisations were often designed for small hill heights $h_{0}$ or in the case where $|f|\ll\omega$.
We also suggested simple adjustments or new parametrisations in the cases where the existing theory falls short. In most cases, only simple adjustments to the standard theory was required, highlighting the relative success of the underlying linear theory, even in regimes considered to be strongly non-linear. It is hoped that these updated parametrisations may be used to more accurately model stress in regimes for which there is currently limited theory available. However, before applying our work to large-scale ocean models, there are some notable limitations that need further consideration.
Firstly, one would need to test how well our scalings and parametrisations hold with more complicated topography, beyond the isolated hill scenario that we studied. For example, as we mentioned in the text, Klymak et al. (2021) found that for steady flows in three-dimensions, the stress varies differently with the buoyancy frequency $N$ depending on whether one has an isolated obstacle, or multiple randomly distributed obstacles. Our assumption that $N$ is uniform should also be relaxed, and as discussed by Lüschow and Von Storch (2024), the choice of $N$ that one uses in a parametrisation can significantly affect the accuracy of the model.
Secondly, there are some specific parameter regimes where more analysis is needed. In particular, more investigation is needed of steady flow regimes where the Froude number $Fr$ has an intermediate value, rather than just the limits $Fr\lesssim 1$ and $Fr\gg 1$ considered here.
Lastly, a critical step in implementing any stress parameterisation is determining where in the vertical the stress should be applied to the flow, which was not considered here. However, previous studies have provided guidance on this point. For mean stress, the propagating wave component is applied where the lee waves break and dissipate (e.g., Booker and Bretherton, 1967; Andrews and Mcintyre, 1976; Xie and Vanneste, 2015), while the non-propagating part acts directly adjoining the hill (e.g., Klymak et al., 2010; Klymak, 2018; Klymak et al., 2021). By contrast, for periodic stress, the force acts in the diffusive boundary layer adjoining the topography as shown in Shakespeare et al. (2021a). However, no previous studies have considered the mixed flow case, in which this picture may differ, and this is clearly an important avenue for further investigation.
Despite these limitations, the present work provides an important step towards a more comprehensive and robust parametrisation of topographic stresses in ocean models.